According to Masterfoods, the company that manufactures M&M’s, 12% of peanut M&M’s are brown, 15% are yellow, 12% are red, 23% are blue, 23% are orange and 15% are green. You randomly select five peanut M&M’s from an extra-large bag of the candies. (Round all probabilities below to four decimal places; i.e. your answer should look like 0.1234, not 0.1234444 or 12.34%.)

Compute the probability that exactly two of the five M&M’s are yellow. Incorrect

Compute the probability that two or three of the five M&M’s are yellow. Incorrect

Compute the probability that at most two of the five M&M’s are yellow. Incorrect

Compute the probability that at least two of the five M&M’s are yellow. Incorrect

If you repeatedly select random samples of five peanut M&M’s, on average how many do you expect to be yellow? (Round your answer to two decimal places.) Incorrect yellow M&M’s

With what standard deviation? (Round your answer to two decimal places.) Incorrect yellow M&M’s
Answer #1

1)

probability that exactly two of the five M&M’s are yellow :

probability = P(X=2)= (nCx)px(1−p)(n-x)    = 0.1382

2)

probability that two or three of the five M&M’s are yellow :

probability = P(2<=X<=3)= x=x1x2     (nCx)px(1−p)(n-x)    = 0.1626

3)

probability that at most two of the five M&M’s are yellow :

probability = P(X<=2)= x=0x     (nCx)px(1−p)(n-x)    = 0.9734

4)

probability = P(X>=2)= 1-P(X<=1)= 1-∑x=0x-1     (nCx)px(1−p)(n-x)    = 0.1648

5)

expect to be yellow =np=0.75

6)

standard deviation σ=√(np(1-p))= 0.80